OpenStudy (mendicant_bias):

(Sturm-Liouville) Problem posted below. My first inclination is to divide through by x to put it into the form of a Cauchy-Euler equation, but I don't know if I'm allowed to do that (that probably sounds silly, oh well).

OpenStudy (mendicant_bias):

\[x^5y''+5x^4y'+\lambda x^3y=0\]

OpenStudy (mendicant_bias):

(Again, first inclination is to divide by x^3 through to put it into the form of a cauchy-euler problem, is this allowed/doable?)

OpenStudy (mendicant_bias):

@dan815

OpenStudy (dan815):

ya

OpenStudy (mendicant_bias):

Alright, cool. So then\[x^2y''+5xy'+\lambda y=0\]

OpenStudy (dan815):

ya this shud have a simple soln

OpenStudy (mendicant_bias):

Yeah, just haven't done cauchy-euler in a while and am trying to remember the general format of it/how it works

OpenStudy (dan815):

y=X^n

OpenStudy (mendicant_bias):

\[y=x^m\]

OpenStudy (mendicant_bias):

Yeah, I gotchu, I think I remember

OpenStudy (mendicant_bias):

\[m(m-1)x^2x^{(m-2)}+5mxy^{(m-1)}+\lambda x^m=0\]

OpenStudy (mendicant_bias):

Simplifying through: \[(m^2-m)x^m+5mx^m+\lambda x^m=0\]That doesn't seem right, one moment.

OpenStudy (mendicant_bias):

Oh nevermind, divide across by x^m, I think? What's the justification for that, can it be shown by bounds or otherwise that x^m is not zero, or is that part of the general conditions on the problem that need to be fulfilled for this to work?

OpenStudy (dan815):

x^m is not zerno unless m is -infinity so

OpenStudy (dan815):

so u can just say X^m is never zero

OpenStudy (dan815):